OpenStudy (anonymous):

a car assembly makes 60 cars in 27 days with the help of 45 artisans. Approximately how many cars can be made in 14days with the help of 15 artisans?

OpenStudy (anonymous):

1 artisan can build \[60 \over {27 * 45}\] cars per day. So, 15 artisans can build \[{{60 * 14 *15} \over {27 * 45}} = 10.37 \] - that is 10 cars - in 2 weeks.

OpenStudy (anonymous):

for 45 artisans: 27days = 60 cars 1 day = 60/25 cars if 45artsans =60/27 cars in a day, then 1 artisan = [(60/27) / 45] cars So 15 artisans in 1 day = [(60/27) / 45] *15 cars In 14 days, 15 artisans will make { [(60/27) / 45] *15 }*14 = 10 cars. Final answer is 10 cars. by Samuel Risbond

OpenStudy (anonymous):

for 45 artisans: 27days = 60 cars 1 day = 60/25 cars if 45artsans =60/27 cars in a day, then 1 artisan = [(60/27) / 45] cars So 15 artisans in 1 day = [(60/27) / 45] *15 cars In 14 days, 15 artisans will make { [(60/27) / 45] *15 }*14 = 10 cars. Final answer is 10 cars. by Samuel Risbond